OpenStudy (anonymous):

Fans & Medals!! I really need help! Erik needs a plumber and must decide between two companies. For a service visit, Plumber A charges $85 to send a plumber, plus $20/h. Plumber B charges $60 to send a plumber, plus $30/h. The system that models this situation is given, where c is the cost of hiring a plumber and h is the number of hours the plumber works. The solution to the system is (2.5, 135). What is the correct interpretation of the solution of this system? c=85+20h c=60+30h

OpenStudy (dtan5457):

c=85+20h c=60+30h IS the right way to interpret this question..are you trying to solve it too?

OpenStudy (anonymous):

Yes

OpenStudy (dtan5457):

Since they are both c= set them both together

OpenStudy (dtan5457):

85+20h=60+30h solve for h

OpenStudy (anonymous):

hold on I have the options=)

OpenStudy (dtan5457):

sure. i have to go in 5 minutes let's make this quick :O

OpenStudy (anonymous):

A. Plumber A will charge more when he works 2.5 h, charging $135. B. Plumber B will charge more when she works 2.5 h, charging $135. C. Plumber A and Plumber B will both charge $135 if they each work 2.5 h. D. Plumber A and Plumber B will both charge $2.50 if they each work 135 h.

OpenStudy (dtan5457):

Knowing the solution is 2.5,135 for the system, it has something to do with BOTH plumbers

OpenStudy (dtan5457):

2.5 hours have both plumbers at the same price ^which of those represent that?

OpenStudy (anonymous):

so im thinking D?

OpenStudy (dtan5457):

"Plumber A and Plumber B will both charge $2.50 if they each work 135 h." i mean really..logically....what plumber will charge $2.50 for 135 hours of work???

OpenStudy (dtan5457):

No it's not D lol

OpenStudy (dtan5457):

It's the other one that has the word both in it ..lol

OpenStudy (anonymous):

OH thanks alot=D

OpenStudy (dtan5457):

no prob.

OpenStudy (dtan5457):

gtg ;p